OpenStudy (anonymous):

Factor: 9x2 - 25y2 (3x + 5y)(3x – 5y) (3x + y)(3x – 25y) (9x + 5y)(x – 5y) prime

OpenStudy (anonymous):

this is the difference of two sqaures the first one is correct

OpenStudy (anonymous):

notice is you expand this the terms in xy cancel out

OpenStudy (anonymous):

wat i dnt understand

OpenStudy (anonymous):

9x^2-25y^2=(3x)^2-(5y)^2 now use a^2-b^2-(a+b)(a-b) and try it out urself please :)

OpenStudy (anonymous):

ok do u want me to show u ?? how i do it??

OpenStudy (anonymous):

i mean the formula is a^2-b^2=(a+b)(a-b) yea show me , how u can get to answer with the formula i've given u.... :)

OpenStudy (anonymous):

davette the two are both Perfect Squares 9x^2 = (3x)^2 and 25y^2 = (5y)^2 so you take the square roots 3x and 5y and place them one in each bracket and pace a + in one bracket and a - inthe other Look out for difference of 2 squares and proceed as above - thats the way I do it!